OpenStudy (anonymous):

Using Law of Sines. a=3 b=7 A=70 I am trying to find the SinB using sin-1(7sin70)/3) but I get domain error.

OpenStudy (jdoe0001):

hmmm can you post a quick screenshot of the material? \(\bf \left[\cfrac{7sin(70^o)}{3}\right]\approx 2.19\qquad recall \implies -1\le sin(\theta)\le 1\)

OpenStudy (anonymous):